OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = –2x2 – 32x – 126? Domain: All Real Numbers Range: y 2 Domain: All Real Numbers Range: y –8 Domain: All Real Numbers Range: y –8 Domain: All Real Numbers Range: y 2

OpenStudy (anonymous):

NatalieLove, If the domain is the set of all real numbers (taken above as it is only option), it is unlikely that there is one single answer for the range such as y=2, or y=-8, and in fact, the first and last option anre the same, as well as the middle two,so perhaps something is wrong here? for example: f(2) = -2(2)^2 -32(2) -126 f(2)= -2(4)-64-126 f(2)=-8-64-126 f(2)=-198 and of course -198 is not in the ranges given. for a range to be a single number(real or integer) the function would have to give the same answer no matter what value is chosen for x.